### v1

# WattsOnAI: Measuring, Analyzing, and Visualizing Energy and Carbon Footprint of AI Workloads

Abstract.
The rapid advancement of AI, particularly large language models (LLMs), has raised significant concerns about the energy use and carbon emissions associated with model training and inference. However, existing tools for measuring and reporting such impacts are often fragmented, lacking systematic metric integration and offering limited support for correlation analysis among them. This paper presents WattsOnAI, a comprehensive software toolkit for the measurement, analysis, and visualization of energy use, power draw, hardware performance, and carbon emissions across AI workloads. By seamlessly integrating with existing AI frameworks, WattsOnAI offers standardized reports and exports fine-grained time-series data to support benchmarking and reproducibility in a lightweight manner. It further enables in-depth correlation analysis between hardware metrics and model performance and thus facilitates bottleneck identification and performance enhancement. By addressing critical limitations in existing tools, WattsOnAI encourages the research community to weigh environmental impact alongside raw performance of AI workloads and advances the shift toward more sustainable ”Green AI” practices. The code is available at https://github.com/SusCom-Lab/WattsOnAI.
copyright: nonedoi: isbn: conference: ; 2025journalyear: ;price:
1. Introduction
Artificial intelligence (AI) is advancing at an extraordinary pace, achieving state-of-the-art performance across domains such as natural language processing (NLP), computer vision (CV), and scientific computing. These breakthroughs are driven by increasingly large models, such as GPT-4, PaLM, and LLaMA, with hundreds of billions or even trillions of parameters trained on massive datasets using large-scale GPU clusters. However, this progress comes with growing environmental and economic costs. Training modern large language models (LLMs) can consume hundreds of megawatt-hours of electricity, with energy demands comparable to the annual usage of a small town (strubell2020energy, ; patterson2021carbon, ). Meanwhile, LLM inference, especially in commercial deployments and operating at web scale, further compounds energy use over time. These trends raise pressing concerns about the sustainability of AI at scale, particularly in the face of global decarbonization goals and resource-constrained infrastructures.
Despite this, the AI research community remains largely focused on improving accuracy, scalability, and latency, with energy and carbon footprint treated as secondary or ignored entirely (schwartz2020green, ; henderson2020towards, ). This omission carries practical, economic, and ethical consequences. On the economic front, a single LLM architecture search or fine-tuning experiment can incur thousands of dollars in cloud computing fees, much of which is attributable to energy-intensive GPU usage (lacoste2019quantifying, ). Environmentally, as AI becomes a larger component of digital infrastructure, its carbon footprint will increasingly come under scrutiny, especially as sustainability reporting frameworks (e.g., the EU’s Corporate Sustainability Reporting Directive (hummel2024overview, )) mandate transparency in operational emissions. This makes it imperative to treat energy/carbon impact as a first-class metric in AI development and deployment.
To address the high energy costs and environmental impact of AI, an important research direction is the development of models capable of predicting energy consumption before task execution. This facilitates more proactive energy management and optimization. While some explorations have been conducted (chen2025much, ), building reliable predictive models for varied AI tasks still requires a solid foundation. A key factor impacting the construction of such models is the need for comprehensive and standardized measurements.
Currently, although some tools are available for measuring and reporting energy and power consumption, they exhibit critical limitations.
Table 1. Feature Comparison of Popular Tools
Tool / Feature Fine metricsa Coarse metricsb GPU power CPU/RAM power Carbon estimation AI task sync. Correlation analysis Multi-dim visualization Complexity& overhead
NVIDIA-SMI (nvidia-smi, ) Limitedc Full Full Limited Limited Limited Limited Limited Low
DCGM (dcgm, ) Full Full Full Limited Limited Limited Limited Limited Low
Nsight (nsys/ncu) (nsight-systems, ; nsight-compute, ) Full Full Full Limited Limited Limited Partial Full High
PyJoules (pyjoules, ) Limited Limited Full Full Limited Full Limited Limited Medium
CarbonTracker (anthony2020carbontracker, ) Limited Limited Full Full Full Full Limited Limited Medium
CodeCarbon (benoit_courty_2024_11171501, ) Limited Limited Full Full Full Full Limited Limited Medium
WattsOnAI (Proposed) Full Full Full Full Full Full Full Full Low
  • a
    Refers to fine-grained metrics (down to the Streaming Multiprocessor (SM) and Tensor Core level), which are included in Appendix A.
  • b
    Refers to coarse-grained metrics, such as those collectible by nvidia-smi (e.g., GPU utilization, GPU memory utilization, temperature, etc.).
  • c
    In the table, ‘Full‘ denotes that the feature is supported, ‘Partial‘ denotes partial support, and ‘Limited‘ denotes that the feature is unsupported.
Fragmented measurement and limited carbon awareness. While various tools, for example, NVIDIA-SMI, DCGM, and Nsight (Section 2.1), exist to collect metrics such as power consumption, energy usage, and hardware utilization for AI tasks, they often overlook carbon emissions and lack integration across these dimensions. This fragmentation hinders the generation of comprehensive, interpretable reports that reflect the true environmental cost of AI workloads. A unified framework that incorporates carbon estimation and supports standardized reporting is urgently needed.
Complexity in usage and operational overhead. Many similar tools are complex to deploy. The setup and configuration of these tools can be intricate. This process often requires specific dependencies or non-trivial integration efforts within diverse AI frameworks, and thus creates a barrier to entry for some users. Additionally, the monitoring process often has notable impacts, including high CPU usage, high memory usage, and significant time consumption. This overhead can interfere with the AI workload’s performance.
Insufficient in multi-dimensional visualization and correlation analysis. While existing tools like CodeCarbon and PyJoules (Section 2.2) support tracking energy or carbon metrics, a major challenge lies in their inadequate visualization of multi-dimensional data such as energy and hardware state. This deficiency indirectly results in a lack of in-depth correlation analysis against hardware characteristics (e.g., GPU type, memory usage) or model performance indicators (e.g., latency, accuracy), hindering the observation of dynamic trends and limiting the fine-grained, data-driven optimization of AI workloads.
By bridging existing research gaps and addressing corresponding challenges, we introduce WattsOnAI and make the following major contributions.
  • Comprehensive Measurement: WattsOnAI unifies real-time monitoring of energy, power, and hardware metrics with carbon emission estimation, offering synchronized tracking throughout AI workload execution. Upon task completion, WattsOnAI generates intuitive post-task reports and exports detailed time-series data, addressing key gaps in metric integration and carbon-aware reporting found in existing tools.
  • Ease of Use and Low Overhead: WattsOnAI is simple to use and seamlessly integrates with various AI workflows; concurrently, it operates with minimal performance overhead (Section 4). This addresses the challenge of a high barrier to entry, avoids impacting AI task performance, and ensures reliable long-term measurement.
  • Multi-Dimensional Visualization and Correlation Analysis: WattsOnAI implements multi-dimensional visualization capabilities, effectively demonstrating how metrics such as energy consumption, carbon emissions, and hardware characteristics evolve over time and relate to one another. Through these tailored visualizations, correlation analysis across classified metrics spanning compute, memory, and communication dimensions is enabled. This bridges a key gap in existing tools by supporting multi-dimensional, workload-aware efficiency analysis.
Ultimately, our aim is not only to provide an effective and comprehensive toolkit for benchmarking energy, power, carbon emissions, and hardware metrics of AI workloads, but also to promote deeper awareness of the environmental costs embedded in AI research and development.
2. Related Work
In this section, we review existing tools (shown in Table 1) capable of measuring energy, power consumption, hardware metrics, and carbon emissions related to AI workloads.
2.1. Hardware Monitoring Interfaces and Libraries
NVIDIA provides Nsight Systems (nsys) (nsight-systems, ) and Nsight Compute (ncu) (nsight-compute, ), which possess deep performance profiling capabilities with fine granularity, effective for identifying computational bottlenecks. However, their focus is on performance debugging, and their high sampling precision (especially ncu) incurs high operational overhead, making them unsuitable as low-interference, long-duration measurement tools.
NVIDIA-SMI (nvidia-smi, ) is a command-line utility provided by NVIDIA for managing and monitoring GPU devices; its overhead is very low, but it only provides coarse-grained metric measurements, such as overall GPU power consumption and utilization. NVIDIA Data Center GPU Manager (DCGM) (dcgm, ) offers finer granularity, capable of acquiring a wide range of relatively fine-grained metrics, including Tensor Core and SM activity status, suitable for large-scale cluster measurements. However, integrating these tools into a unified measurement framework precisely synchronized with the AI task execution process and including energy consumption and carbon emission estimation requires researchers to exert additional effort for secondary development and metric integration (such as Intel RAPL (rapl, )).
Figure 1. The architecture of WattsOnAI.
2.2. Energy Measurement Frameworks and Tools
PyJoules (pyjoules, ) is a Python library that utilizes interfaces like Intel RAPL (rapl, ) and NVIDIA NVML (nvml, ) to retrieve hardware energy consumption data. CodeCarbon (benoit_courty_2024_11171501, ) is a relatively mature tool specifically focused on estimating the carbon emissions produced during code execution. It operates by measuring the energy consumption of key hardware components (such as CPU, GPU, and RAM) and then applying location-based carbon intensity factors to convert this energy consumption into a carbon footprint. CarbonTracker (anthony2020carbontracker, ) is also a tool with similar functionality to CodeCarbon. Besides these, there are also system-level measurement and cloud platform tools, but their granularity is generally coarser.
Existing energy and carbon measurement tools exhibit several shortcomings. Firstly, these tools typically report only total or average values, which restricts their capability for deep correlation analysis and makes it difficult to attribute specific energy or power consumption changes to underlying hardware events during AI tasks. Secondly, their visualization capabilities tend to concentrate on single metrics, such as carbon emissions, often lacking the interactive, multi-dimensional perspectives needed to easily interpret complex system behaviors.
3. Architecture and Design
Fig. 1 shows the architecture of WattsOnAI, which is fundamentally divided into i) source interface layer, ii) data processing and persistance layer, and iii) demonstration layer. Through this three-layer architecture, WattsOnAI systematically measures, analyses and visualizes energy consumption, power consumption, carbon emissions and other relevant hardware metrics throughout the execution of AI tasks.
3.1. Source Interface Layer
As mentioned in the previous section, metric data collection primarily relies on interfaces provided by the hardware. Based on user configuration, WattsOnAI dynamically selects the metrics to sample and can support up to 26 distinct metrics. These include GPU power consumption, SM active, among others.
Furthermore, a key design goal is to achieve the highest sampling frequency permitted by the underlying hardware interfaces. This high temporal resolution allows researchers to observe subtle variations in energy consumption and power draw across different fine-grained stages within AI task execution.
Given that different hardware components expose metrics through distinct interfaces, and even a single component like an NVIDIA GPU might offer data via multiple mechanisms (e.g., NVML/SMI and DCGM), we employ a parallelized collection strategy. This is handled within WattsOnAI by our MetricsCollector class, which is responsible for the statistical metric collection. Where appropriate (primarily for GPU metrics), this allows WattsOnAI to query interfaces like those underlying SMI and DCGM concurrently. This approach maximizes the achievable sampling rate to approximately 0.1s-0.2s per sample, a frequency no lower than that of pynvml (pynvml, ).
3.2. Data Processing and Persistence Layer
After collected from the hardware, the raw data go through the following two-stage converting processes.
Stage-1: Normalization involves converting cumulative energy metrics (e.g., Joules) obtained for CPU and DRAM via interfaces into average power consumption (e.g., Watts) over the sampling interval. This unit regularization simplifies subsequent analysis and offers a more intuitive view of resource usage.
Stage-2: Energy consumption data is converted into estimated carbon emissions. This conversion is critical for meaningful environmental impact reporting, as grid carbon intensity factors vary significantly across both geography and time. As demonstrated by article (wu2025carbonedge, ), substantial real-time carbon intensity differences exist even at mesoscales (tens to hundreds of km), with simultaneous variations reaching factors of 7.9× (Western US) and 19.5× (Central Europe) between adjacent areas. Furthermore, significant temporal fluctuations occur, including intraday and seasonal variations. This evidence strongly emphasizes the dynamic nature of carbon intensity.
This conversion process raises two key questions. The first concerns the distinction between embodied and operational carbon (li2024carbon, ).
  • Embodied carbon encompasses the emissions from manufacturing, transport, etc., representing a one-time upfront cost associated with the hardware itself.
  • Operational carbon stems from the energy consumed during the hardware’s use phase.
While embodied carbon is crucial for comparing different hardware solutions or for full Life Cycle Assessments (LCA), WattsOnAI’s primary focus is on quantifying the direct impact of specific AI tasks (training, inference, etc.) during their execution. Consequently, we deliberately exclude embodied carbon and measure only the operational carbon generated during task execution, using the following rule:
(1) $Carbon=Energy\times Intensity.$
The second consideration is the type of carbon intensity factor used. Average intensity reflects the grid’s overall generation mix and is suited for broad estimations. Marginal intensity, however, represents the emissions from the power source(s) activated to meet the additional load imposed by the task. Since this marginal generation often comes from more carbon-intensive sources (like fossil fuel peaker plants), its intensity is typically higher and more accurately reflects the direct environmental consequence of running the AI task. To provide a clearer picture of the incremental impact and environmental cost, WattsOnAI uses marginal carbon intensity for its estimations, with the necessary intensity factors being retrieved from services like Electricity Maps (electricity_maps, ) and WattTime (Watttime, ) based on geographical coordinates.
Finally, all processed and normalized data, including these carbon estimates, are strictly time-aligned within the data series before being persistently stored in files or a database.
3.3. Demonstration Layer
The demonstration layer retrieves the processed and stored data, potentially applying final transformations for display, and presents it to the user. The goal is to provide multi-faceted data visualization options, enabling intuitive and comprehensive analysis. Key demonstration methods include:
  • Execution Report: After the task finishes, a summary report is generated.
  • Tabular Data: Provides detailed, time-stamped metric information in a table format, suitable for export (e.g., CSV) or direct inspection.
  • Plotting: Users can generate visualizations, typically line graphs, from the stored time-series data to observe the behavior of metrics over the task’s duration.
  • Real-time Monitoring Dashboard: We also integrate WattsOnAI with visualization tools (e.g., Grafana (Grafana, )) for live monitoring during task execution.
These multiple visualization approaches (referring to Appendix B for examples) combine to offer a more detailed and comprehensive view of the collected data.
4. Case Study
from WattsOnAI import monitor
try:
monitor.start(sampling_interval = 0.1, …)
# Llama2-7b inference (sleep 15s between Prefill and Decode)
finally:
monitor.stop()
Utilizing WattsOnAI as demonstrated in the code snippet above, we conducted an LLM inference experiment using the Llama2-7b model on a server with NVIDIA A800 GPU, and sampled All Metrics (see Appendix A for the full list) at a sampling interval of $0.1$ second. With the help of WattsOnAI, this experiment yields findings with the following insights and implications.
4.1. Correlation Analysis on Phase Dynamics
Figure 2. Power and frequency during inference phase.
For inference tasks, the LLM sequentially goes through the Prefill phase and Decode phase in generating the output (tokens). Observations from Fig. 2 indicate that the Prefill phase has significantly higher peak power consumption (exceeding the Decode phase by nearly 90W) and exhibits larger power fluctuations. Consequently, the Prefill phase is the primary contributor to peak power demand, making optimization of its computational efficiency (e.g., using techniques like FlashAttention) crucial for reducing overall peak power. GPU frequency patterns mirror this, with the Prefill phase operating at lower and more variable frequencies, while the Decode phase sustains higher and more stable frequencies.
Figure 3. Shifting bottlenecks between phases.
The two phases also differ in their primary performance bottlenecks, as shown in Fig. 3. The Prefill phase is predominantly compute-bound, whereas the Decode phase is mainly memory-bound. The underlying reasons stem from their operational nature: Prefill compute-intensively processes the full input in parallel, whereas Decode sequentially generates tokens, frequently accessing the large KV cache for smaller computations to predict the next token.
Figure 4. Occupancy and active during decode phase.
Further analysis of GPU utilization during the inference process (Fig. 4) highlights a strong inverse correlation between SM occupancy and Tensor active percentages. A reasonable explanation is that the GPU switches its main effort during inference: when it is doing heavy math, the specialized Tensor Cores are busy and SMs mostly assist; when it is doing other general tasks, the SMs are busy ones and the Tensor Cores are quieter. As previously analyzed and validated in Fig. 4, the Decode phase is not compute-intensive and thus the Tensor active often remains low during this phase.
Figure 5. Carbon emissions from the same AI task in two neighboring provinces of Canada, respectively.
4.2. Analysis on Carbon Emission Estimation
In the aforementioned experiment, the report generated by WattsOnAI indicated that the task consumed 7485 Joules (2.08 $Watts\cdot{hour}$) of energy. However, as we have mentioned, such energy consumption figure does not fully reflect the task’s actual environmental impact, which motivated our introduction of the carbon emissions metric.
To demonstrate the decisive role of geographical location, we performed a comparative analysis: assuming that this task was executed in the Canadian provinces of Saskatchewan and Manitoba respectively, Fig. 5 shows that the identical AI task could yield a difference in generated carbon emissions, 1.03g vs 0.07g, approaching 15-fold between these neighboring provinces.
This clearly indicates that geographical location is a critical factor when assessing the environmental footprint of AI tasks. This finding aligns with research such as (nguyen2024towards, ), who also demonstrate significant variations in operational carbon emissions for identical LLM tasks when executed in regions with differing grid carbon intensities, such as Québec, California, and the US PACE territories. Indeed, converting energy consumption to carbon emissions using location-specific carbon intensity factors, as employed in our analysis and detailed by (nguyen2024towards, ), is crucial for accurate environmental assessments.
4.3. Overhead Caused by WattsOnAI
We evaluated the system overhead of WattsOnAI using pidstat across two control groups. The first group measured the inherent overhead of WattsOnAI by running it on an otherwise idle system with a 0.1-second sampling interval. The second group replicated the experiment from Section 4, but with the modification that we removed the 15-second sleep time between the prefill and decode phases. In this setup, we evaluated the impact of WattsOnAI with sampling intervals of 0.1, 0.5, 1, and 5 seconds. The results are presented in Table 2.
Table 2. Overhead Caused by WattsOnAI (WOA)
Overhead Elapsed time Time Overhead Time /sample CPU util. Memory cons.
Baseline 15s 0% 121.21MB
WOA(0.1) only 15.05s 0.33%  $\uparrow$ 0.37ms 2.23% 121.29MB
Exp w/o WOA 29.15s 99.1% 1411.37MB
Exp w WOA(5) 29.39s 0.82%  $\uparrow$ 38.51ms 97.21% 1388.59MB
Exp w WOA(1) 29.53s 1.30%  $\uparrow$ 12.46ms 97.67% 1398.04MB
Exp w WOA(0.5) 29.71s 1.92%  $\uparrow$ 9.28ms 98.24% 1398.84MB
Exp w WOA(0.1) 30.54s 4.77%  $\uparrow$ 7.2ms 98.97% 1403.62MB
The data in the table indicates that the overhead caused by WattsOnAI is low. Besides, it can be seen that the higher the sampling frequency, the higher the overhead, but the time consumption per sample is smaller. Regarding CPU utilization, the highest usage is observed when WattsOnAI is absent, as the tool itself consumes a small portion of CPU resources. Finally, the differences in memory consumption are negligible and can be considered measurement noise.
5. Conclusion and future work
Aiming to foster greater awareness of energy consumption in AI research, we developed a software toolkit WattsOnAI, which enables measuring, analyzing, and visualizing energy and carbon footprint of AI workloads. WattsOnAI also forms a solid foundation for AI energy predicting models. Future work will focus on achieving task-level granularity (e.g., attributing energy consumption to individual processes running concurrently on shared hardware) and extending compatibility beyond NVIDIA GPUs.
We hope this work contributes meaningfully to the advancement of Green AI principles, inspires further research into sustainable AI practices, and supports efforts to reduce the carbon footprint of modern AI systems.
Appendix A Metric Explanation
Table 3 details the collected GPU performance metrics, categorized by their primary function.
Table 3. GPU Performance Metrics
Section Metric and Description
Energy Section power.draw [W]: Current real-time power consumption of the GPU, in Watts (W). temperature.gpu: Current temperature of the main GPU core. cpu_power: Current real-time power consumption of the CPU, in Watts (W). dram_power: Current real-time power consumption of the DRAM, in Watts (W).
Compute Section utilization.gpu [%]: Percentage of time over the past sample period during which one or more kernels were executing on the GPU’s Streaming Multiprocessors (SMs). sm_active: Percentage of time and quantity the Streaming Multiprocessors (SMs) were active (executing instructions). sm_occupancy: Ratio of active warps on a Streaming Multiprocessor (SM) to the maximum number of warps supported by the SM. tensor_active: Percentage of time and quantity the Tensor Cores (used for accelerating AI computations) were active. fp64_active: Percentage of time the GPU’s double-precision (FP64) units were active. fp32_active: Percentage of time the GPU’s single-precision (FP32) units were active. fp16_active: Percentage of time the GPU’s half-precision (FP16) units were active. clocks.current.graphics [MHz]: Current clock frequency of the GPU’s graphics/shader cores, in Megahertz (MHz). clocks.current.sm [MHz]: Current clock frequency of the GPU’s Streaming Multiprocessors (SMs), in Megahertz (MHz).
Memory Section utilization.memory [%]: Percentage of time over the past sample period during which the GPU’s memory interface was busy. dram_active: The proportion of cycles the interface was actively transferring data, reflecting bandwidth usage efficiency (Analogy: Consider utilization.memory as the time a kitchen pantry door is open, and dram_active as the time a chef’s hands are actually carrying ingredients from it. The door might be open longer than ingredients are being moved.). usage.memory [%]: Percentage of total available GPU memory that is currently allocated or used. temperature.memory: Current temperature of the GPU memory modules. clocks.current.memory [MHz]: Current clock frequency of the GPU memory, in Megahertz (MHz).
Communication Section pcie.link.gen.current: Current generation of the PCIe link, determining the maximum theoretical transfer rate. pcie.link.width.current: Current number of active PCIe lanes used by the link. pcie_tx_bytes: Total number of bytes transmitted from the GPU to the host via the PCIe bus. pcie_rx_bytes: Total number of bytes received by the GPU from the host via the PCIe bus. nvlink_tx_bytes: Total number of bytes transmitted via the NVLink high-speed interconnect. nvlink_rx_bytes: Total number of bytes received via the NVLink high-speed interconnect.
System Section cpu_usage: Percentage of the host system’s CPU utilization. dram_usage: Percentage of the host system’s main memory (RAM/DRAM) currently in use.
Appendix B Demonstration Examples
Tables 45, Figs. 6, and 7 present examples of the four data demonstration approaches.
Table 4. Execution Report
Overall Performance Metric Value
Total Time [s] 43.00
CPU Energy [J] 6348.88
DRAM Energy [J] 388.74
GPU 0 Energy [J] 8918.24
Total Energy [J] 15655.86
Carbon Emissions [kg CO2eq] 0.0020
Component Metric Avg Max Min Mode
CPU Usage [%] 3.67 6.20 2.20 4.00
Power [W] 147.91 160.26 123.73 123.73
DRAM Usage [%] 4.10 4.10 3.90 4.10
Power [W] 9.05 9.43 8.77 9.06
GPU 0 Detailed Statistics (NVIDIA A800 80GB PCIe)
Category Metric Avg Max Min Mode
Energy Power Draw [W] 204.12 315.58 62.89 65.86
GPU Temp. [°C] 43.09 51.00 32.00 38.00
Compute GPU Utilization [%] 63.55 100.00 0.00 100.00
Graphics Clock [MHz] 1354.43 1410.00 1125.00 1410.00
SM Clock [MHz] 1354.43 1410.00 1125.00 1410.00
SM Active [%] 59.91 100.00 0.00 100.00
SM Occupancy [%] 20.16 92.30 0.00 0.00
Tensor Active [%] 19.91 93.70 0.00 0.00
FP64 Active [%] 0.00 0.00 0.00 0.00
FP32 Active [%] 1.45 31.90 0.00 0.00
FP16 Active [%] 0.13 5.10 0.00 0.00
Memory Mem. Utilization [%] 27.41 54.00 0.00 0.00
Mem. Temp. [°C] 45.91 55.00 35.00 41.00
Mem. Clock [MHz] 1512.00 1512.00 1512.00 1512.00
Mem. Usage Total [%] 88.94 94.56 17.23 90.02
DRAM Active (Cycles) [%] 18.40 94.60 0.00 0.00
Communication PCIe Link Gen 4.00 4.00 4.00 4.00
PCIe Width 16.00 16.00 16.00 16.00
PCIe TX [GB/s] 0.11 0.14 0.07 0.10
PCIe RX [GB/s] 0.05 0.13 0.04 0.05
NVLink TX [GB/s] 0.00 0.00 0.00 0.00
NVLink RX [GB/s] 0.00 0.00 0.00 0.00
Top Positively and Negatively Correlated Metric Pairs
Metric A Metric B Coeff. Metric A Metric B Coeff.
SM Clock [MHz] Graphics Clock [MHz] 1.000 SM Clock [MHz] Tensor Active [%] -0.844
GPU Utilization [%] Power Draw [W] 0.966 Tensor Active [%] Graphics Clock [MHz] -0.844
SM Active [%] Power Draw [W] 0.948 SM Active [%] PCIe TX [GB/s] -0.807
GPU Utilization [%] SM Active [%] 0.938 PCIe TX [GB/s] Power Draw [W] -0.768
Mem. Temp. [°C] GPU Temp. [°C] 0.924 GPU Utilization [%] PCIe TX [GB/s] -0.650
Table 5. Tabular Data
Figure 6. Plotting of GPU activities.
Figure 7. Real-time monitoring dashboard.

### v2

@@
-WattsOnAI: Measuring, Analyzing, and Visualizing Energy and Carbon Footprint of AI Workloads
+AIMeter: Measuring, Analyzing, and Visualizing Energy and Carbon Footprint of AI Workloads
 Abstract.
@@
 However, existing tools for measuring and reporting such impacts are often fragmented, lacking systematic metric integration and offering limited support for correlation analysis among them.
-This paper presents WattsOnAI, a comprehensive software toolkit for the measurement, analysis, and visualization of energy use, power draw, hardware performance, and carbon emissions across AI workloads.
-By seamlessly integrating with existing AI frameworks, WattsOnAI offers standardized reports and exports fine-grained time-series data to support benchmarking and reproducibility in a lightweight manner.
+This paper presents AIMeter, a comprehensive software toolkit for the measurement, analysis, and visualization of energy use, power draw, hardware performance, and carbon emissions across AI workloads.
+By seamlessly integrating with existing AI frameworks, AIMeter offers standardized reports and exports fine-grained time-series data to support benchmarking and reproducibility in a lightweight manner.
 It further enables in-depth correlation analysis between hardware metrics and model performance and thus facilitates bottleneck identification and performance enhancement.
-By addressing critical limitations in existing tools, WattsOnAI encourages the research community to weigh environmental impact alongside raw performance of AI workloads and advances the shift toward more sustainable ”Green AI” practices.
-The code is available at https://github.com/SusCom-Lab/WattsOnAI.
-copyright: nonedoi: isbn: conference: ; 2025journalyear: ;price:
+By addressing critical limitations in existing tools, AIMeter encourages the research community to weigh environmental impact alongside raw performance of AI workloads and advances the shift toward more sustainable ”Green AI” practices.
+The code is available at https://github.com/SusCom-Lab/AIMeter.
+copyright: noneconference: ; 2025
 1. Introduction
@@
 CodeCarbon (benoit_courty_2024_11171501, ) Limited Limited Full Full Full Full Limited Limited Medium
-WattsOnAI (Proposed) Full Full Full Full Full Full Full Full Low
+AIMeter (Proposed) Full Full Full Full Full Full Full Full Low
   • a
@@
 This deficiency indirectly results in a lack of in-depth correlation analysis against hardware characteristics (e.g., GPU type, memory usage) or model performance indicators (e.g., latency, accuracy), hindering the observation of dynamic trends and limiting the fine-grained, data-driven optimization of AI workloads.
-By bridging existing research gaps and addressing corresponding challenges, we introduce WattsOnAI and make the following major contributions.
-  • Comprehensive Measurement: WattsOnAI unifies real-time monitoring of energy, power, and hardware metrics with carbon emission estimation, offering synchronized tracking throughout AI workload execution.
-Upon task completion, WattsOnAI generates intuitive post-task reports and exports detailed time-series data, addressing key gaps in metric integration and carbon-aware reporting found in existing tools.
-  • Ease of Use and Low Overhead: WattsOnAI is simple to use and seamlessly integrates with various AI workflows; concurrently, it operates with minimal performance overhead (Section 4).
+By bridging existing research gaps and addressing corresponding challenges, we introduce AIMeter and make the following major contributions.
+  • Comprehensive Measurement: AIMeter unifies real-time monitoring of energy, power, and hardware metrics with carbon emission estimation, offering synchronized tracking throughout AI workload execution.
+Upon task completion, AIMeter generates intuitive post-task reports and exports detailed time-series data, addressing key gaps in metric integration and carbon-aware reporting found in existing tools.
+  • Ease of Use and Low Overhead: AIMeter is simple to use and seamlessly integrates with various AI workflows; concurrently, it operates with minimal performance overhead (Section 4).
 This addresses the challenge of a high barrier to entry, avoids impacting AI task performance, and ensures reliable long-term measurement.
-  • Multi-Dimensional Visualization and Correlation Analysis: WattsOnAI implements multi-dimensional visualization capabilities, effectively demonstrating how metrics such as energy consumption, carbon emissions, and hardware characteristics evolve over time and relate to one another.
+  • Multi-Dimensional Visualization and Correlation Analysis: AIMeter implements multi-dimensional visualization capabilities, effectively demonstrating how metrics such as energy consumption, carbon emissions, and hardware characteristics evolve over time and relate to one another.
 Through these tailored visualizations, correlation analysis across classified metrics spanning compute, memory, and communication dimensions is enabled.
@@
 However, integrating these tools into a unified measurement framework precisely synchronized with the AI task execution process and including energy consumption and carbon emission estimation requires researchers to exert additional effort for secondary development and metric integration (such as Intel RAPL (rapl, )).
-Figure 1. The architecture of WattsOnAI.
+Figure 1. The architecture of AIMeter.
 2.2. Energy Measurement Frameworks and Tools
@@
 Fig.
-1 shows the architecture of WattsOnAI, which is fundamentally divided into i) source interface layer, ii) data processing and persistance layer, and iii) demonstration layer.
-Through this three-layer architecture, WattsOnAI systematically measures, analyses and visualizes energy consumption, power consumption, carbon emissions and other relevant hardware metrics throughout the execution of AI tasks.
+1 shows the architecture of AIMeter, which is fundamentally divided into i) source interface layer, ii) data processing and persistence layer, and iii) demonstration layer.
+Through this three-layer architecture, AIMeter systematically measures, analyzes and visualizes energy consumption, power consumption, carbon emissions and other relevant hardware metrics throughout the execution of AI tasks.
 3.1. Source Interface Layer
 As mentioned in the previous section, metric data collection primarily relies on interfaces provided by the hardware.
-Based on user configuration, WattsOnAI dynamically selects the metrics to sample and can support up to 26 distinct metrics.
+Based on user configuration, AIMeter dynamically selects the metrics to sample and can support up to 26 distinct metrics.
 These include GPU power consumption, SM active, among others.
@@
 Given that different hardware components expose metrics through distinct interfaces, and even a single component like an NVIDIA GPU might offer data via multiple mechanisms (e.g., NVML/SMI and DCGM), we employ a parallelized collection strategy.
-This is handled within WattsOnAI by our MetricsCollector class, which is responsible for the statistical metric collection.
-Where appropriate (primarily for GPU metrics), this allows WattsOnAI to query interfaces like those underlying SMI and DCGM concurrently.
+This is handled within AIMeter by our MetricsCollector class, which is responsible for the statistical metric collection.
+Where appropriate (primarily for GPU metrics), this allows AIMeter to query interfaces like those underlying SMI and DCGM concurrently.
 This approach maximizes the achievable sampling rate to approximately 0.1s-0.2s per sample, a frequency no lower than that of pynvml (pynvml, ).
@@
   • Operational carbon stems from the energy consumed during the hardware’s use phase.
-While embodied carbon is crucial for comparing different hardware solutions or for full Life Cycle Assessments (LCA), WattsOnAI’s primary focus is on quantifying the direct impact of specific AI tasks (training, inference, etc.) during their execution.
+While embodied carbon is crucial for comparing different hardware solutions or for full Life Cycle Assessments (LCA), AIMeter’s primary focus is on quantifying the direct impact of specific AI tasks (training, inference, etc.) during their execution.
 Consequently, we deliberately exclude embodied carbon and measure only the operational carbon generated during task execution, using the following rule:
@@
 Since this marginal generation often comes from more carbon-intensive sources (like fossil fuel peaker plants), its intensity is typically higher and more accurately reflects the direct environmental consequence of running the AI task.
-To provide a clearer picture of the incremental impact and environmental cost, WattsOnAI uses marginal carbon intensity for its estimations, with the necessary intensity factors being retrieved from services like Electricity Maps (electricity_maps, ) and WattTime (Watttime, ) based on geographical coordinates.
+To provide a clearer picture of the incremental impact and environmental cost, AIMeter uses marginal carbon intensity for its estimations, with the necessary intensity factors being retrieved from services like Electricity Maps (electricity_maps, ) and WattTime (Watttime, ) based on geographical coordinates.
 Finally, all processed and normalized data, including these carbon estimates, are strictly time-aligned within the data series before being persistently stored in files or a database.
@@
   • Plotting: Users can generate visualizations, typically line graphs, from the stored time-series data to observe the behavior of metrics over the task’s duration.
-  • Real-time Monitoring Dashboard: We also integrate WattsOnAI with visualization tools (e.g., Grafana (Grafana, )) for live monitoring during task execution.
+  • Real-time Monitoring Dashboard: We also integrate AIMeter with visualization tools (e.g., Grafana (Grafana, )) for live monitoring during task execution.
 These multiple visualization approaches (referring to Appendix B for examples) combine to offer a more detailed and comprehensive view of the collected data.
 4. Case Study
-from WattsOnAI import monitor
+from AIMeter import monitor
 try:
@@
 monitor.stop()
-Utilizing WattsOnAI as demonstrated in the code snippet above, we conducted an LLM inference experiment using the Llama2-7b model on a server with NVIDIA A800 GPU, and sampled All Metrics (see Appendix A for the full list) at a sampling interval of $0.1$ second.
-With the help of WattsOnAI, this experiment yields findings with the following insights and implications.
+Utilizing AIMeter as demonstrated in the code snippet above, we conducted an LLM inference experiment using the Llama2-7b model on a server with NVIDIA A800 GPU, and sampled All Metrics (see Appendix A for the full list) at a sampling interval of $0.1$ second.
+With the help of AIMeter, this experiment yields findings with the following insights and implications.
 4.1. Correlation Analysis on Phase Dynamics
@@
 4.2. Analysis on Carbon Emission Estimation
-In the aforementioned experiment, the report generated by WattsOnAI indicated that the task consumed 7485 Joules (2.08 $Watts\cdot{hour}$) of energy.
+In the aforementioned experiment, the report generated by AIMeter indicated that the task consumed 7485 Joules (2.08 $Watts\cdot{hour}$) of energy.
 However, as we have mentioned, such energy consumption figure does not fully reflect the task’s actual environmental impact, which motivated our introduction of the carbon emissions metric.
@@
 Indeed, converting energy consumption to carbon emissions using location-specific carbon intensity factors, as employed in our analysis and detailed by (nguyen2024towards, ), is crucial for accurate environmental assessments.
-4.3. Overhead Caused by WattsOnAI
-We evaluated the system overhead of WattsOnAI using pidstat across two control groups.
-The first group measured the inherent overhead of WattsOnAI by running it on an otherwise idle system with a 0.1-second sampling interval.
+4.3. Overhead Caused by AIMeter
+We evaluated the system overhead of AIMeter using pidstat across two control groups.
+The first group measured the inherent overhead of AIMeter by running it on an otherwise idle system with a 0.1-second sampling interval.
 The second group replicated the experiment from Section 4, but with the modification that we removed the 15-second sleep time between the prefill and decode phases.
-In this setup, we evaluated the impact of WattsOnAI with sampling intervals of 0.1, 0.5, 1, and 5 seconds.
+In this setup, we evaluated the impact of AIMeter with sampling intervals of 0.1, 0.5, 1, and 5 seconds.
 The results are presented in Table 2.
-Table 2. Overhead Caused by WattsOnAI (WOA)
+Table 2. Overhead Caused by AIMeter (AIM)
 Overhead Elapsed time Time Overhead Time /sample CPU util. Memory cons.
 Baseline 15s 0% 121.21MB
-WOA(0.1) only 15.05s 0.33%  $\uparrow$ 0.37ms 2.23% 121.29MB
-Exp w/o WOA 29.15s 99.1% 1411.37MB
-Exp w WOA(5) 29.39s 0.82%  $\uparrow$ 38.51ms 97.21% 1388.59MB
-Exp w WOA(1) 29.53s 1.30%  $\uparrow$ 12.46ms 97.67% 1398.04MB
-Exp w WOA(0.5) 29.71s 1.92%  $\uparrow$ 9.28ms 98.24% 1398.84MB
-Exp w WOA(0.1) 30.54s 4.77%  $\uparrow$ 7.2ms 98.97% 1403.62MB
-The data in the table indicates that the overhead caused by WattsOnAI is low.
+AIM(0.1) only 15.05s 0.33%  $\uparrow$ 0.37ms 2.23% 121.29MB
+Exp w/o AIM 29.15s 99.1% 1411.37MB
+Exp w AIM(5) 29.39s 0.82%  $\uparrow$ 38.51ms 97.21% 1388.59MB
+Exp w AIM(1) 29.53s 1.30%  $\uparrow$ 12.46ms 97.67% 1398.04MB
+Exp w AIM(0.5) 29.71s 1.92%  $\uparrow$ 9.28ms 98.24% 1398.84MB
+Exp w AIM(0.1) 30.54s 4.77%  $\uparrow$ 7.2ms 98.97% 1403.62MB
+The data in the table indicates that the overhead caused by AIMeter is low.
 Besides, it can be seen that the higher the sampling frequency, the higher the overhead, but the time consumption per sample is smaller.
-Regarding CPU utilization, the highest usage is observed when WattsOnAI is absent, as the tool itself consumes a small portion of CPU resources.
+Regarding CPU utilization, the highest usage is observed when AIMeter is absent, as the tool itself consumes a small portion of CPU resources.
 Finally, the differences in memory consumption are negligible and can be considered measurement noise.
 5. Conclusion and future work
-Aiming to foster greater awareness of energy consumption in AI research, we developed a software toolkit WattsOnAI, which enables measuring, analyzing, and visualizing energy and carbon footprint of AI workloads.
-WattsOnAI also forms a solid foundation for AI energy predicting models.
+Aiming to foster greater awareness of energy consumption in AI research, we developed a software toolkit AIMeter, which enables measuring, analyzing, and visualizing energy and carbon footprint of AI workloads.
+AIMeter also forms a solid foundation for AI energy predicting models.
 Future work will focus on achieving task-level granularity (e.g., attributing energy consumption to individual processes running concurrently on shared hardware) and extending compatibility beyond NVIDIA GPUs.